OpenStudy (anonymous):

Write an equation of the cosine function with amplitude 2 and period 6pi.

OpenStudy (tkhunny):

\(y = \cos(x)\) Amplitude: \(1\) and Period: \(2\pi\) \(y = 2\cos(x)\) Amplitude: \(2\) and Period: \(2\pi\) \(y = 2\cos((what?)x)\) Amplitude: \(2\) and Period: \(6\pi\)

OpenStudy (anonymous):

1/3?

OpenStudy (tkhunny):

\(\dfrac{2\pi}{1/3} = 6\pi\). Perfect. Why do yo doubt?